# Developers’ Perspectives on Software Licensing: Current Practices, Challenges, and Tools

### v1

@@ 4.2.3. Understanding licenses
In interviews, we sought to clarify how developers acquired an understanding of software licenses. $RI_{3}$, who had indicated that they had received informal training (see Table 1), described a “bootcamp” for licensing consisting of “three sessions of 45 minutes.” Some interviewees described how they had gained an understanding of common licenses over the years, such as $RI_{5}$, who noted, “I generally know that if it’s MIT or Apache, I’m good to go… and those are the ones that I encounter most regularly.” $RI_{6}$ had a similar familiarity with their most-used licenses but acknowledged that this did not extend to licenses they had less experience with: “I’m familiar with… the [Creative Commons] sharealike, the MIT license… But I wouldn’t say that I’m comfortable knowing what all the most common ones mean. I know the ones that I know because they are the ones that I use.”
Interviews also revealed how developers learned about unfamiliar licenses encountered during development. For example, $RI_{7}$ described “[searching] mainly on the web, and looking at other cases. And also the basic Wikipedia…. [I once used] ChatGPT… [for] interpreting the rules that are written [and] making sure that I understood well.” Developers may also look for key terms in software licenses that pertain to them, as explained by $RI_{5}$: “Mostly what I’m looking for is, what are the restrictions on how to use it, and then, what are the particular attribution requirements or restrictions for it.” The terms of interest can vary depending on the use case or organizational policies, such as in the case of $RI_{4}$, who noted that “as far as [my organization] goes, there is a strict policy for patent licensing and getting licenses for IP strategy.”
Finding 9: Developers report best understanding licenses they use most frequently, but may search online or use resources to understand other licenses, paying attention to terms that they believe most affect them.
@@ 4.3. RQ3: Licensing Tools
To understand if and how developers use tools to assist with software licensing tasks, we asked participants to describe their licensing tool use. Here, we refer to a “licensing tool” as any software that can assist with one or more software licensing tasks.
@@ 4.3.1. Tool adoption
Interview participants elaborated, suggesting that the frequency with which licensing tasks are performed may be a factor in low tool use. $RI_{5}$ stated, “I’m not aware of tools [because of] the fact that [licensing is] not something that I have to deal with every day…” Similarly, $RI_{6}$ was unfamiliar with tools, but mused that someone who engaged with licensing more might be more familiar: “I am less aware [of licensing tools]… [but] I would be completely unsurprised if there are established open source tools to assist in this [that] I am not personally aware of, but the people that work for me and do this when it is an important task very well may [know of them].”
Finding 10: Adoption of software licensing tools was low among respondents, who suggested that this may be due to the relatively low frequency at which licensing tasks are performed.
@@ 4.3.2. Tool benefits and shortcomings
Respondents also indicated that tools could help them to choose licenses to assign to their software. In their survey response, $RI_{7}$ indicated that they benefited from tools that could give “an overview of all licensing options [and support] the selection of the right option.” They also noted a capability of “[g]enerating terms that I can copy/paste,” though it is not clear whether this referred to providing the text of existing licenses or creating new licenses. In interviews, $RI_{5}$ and $RI_{7}$ both referred to using choosealicense.com (GitHub, Inc., [n. d.]) to assist them when assigning licenses to software. Specifically, they appreciated its ability to “answer your questions and then recommend the best license,” as $RI_{7}$ put it.
Finding 11: Developers highlighted the capability of tools to assist with identifying licenses in dependencies, highlight those which may cause licensing issues, and choose appropriate licenses for their software as being particularly beneficial.
We also asked respondents who had used licensing tools to describe any particular difficulties or shortcomings that they had encountered. Two respondents indicated that they had encountered issues with licensing tools being cumbersome or difficult to use, with $R_{43}$ describing usability as “poor.” Others noted that tools were not able to account for user feedback: in their survey response, $RI_{7}$ called current tooling “[n]ot interactive enough,” and noted that it could be improved with an “[i]nteractive decision tree” or a “[r]eliable chat bot.” Other concerns included the clarity of tools’ outputs, such as $R_{43}$’s observation that the “[a]ctual effect of a license is obscure[d] to developers,” and $R_{37}$’s opinion that tools “[m]ay abstract away finer details.”
Finding 12: In the views of respondents, licensing tools are hindered by low usability, low interactivity, and a lack of clarity and depth to analysis.
@@ 4.3.3. Areas requiring better tool support
Given the low reported tool adoption, in interviews, we investigated which licensing tasks respondents believed could most benefit from better tool support.
In interviews, respondents noted that ways to better explain licenses and their terms would be beneficial. To illustrate this, $RI_{5}$ cited how “GitHub [has a] drop-down list when you choose a license, but it doesn’t tell you what [the licenses] mean.” $RI_{3}$ echoed this sentiment, elaborating that interpreting the terms of software licenses was a challenge: “[T]here is a gap between what the license gives you as $\dots$ content and [how] you can interpret it $\dots$ [You] can understand the text, but if you need to learn how to interpret that, that would be nice, a tool that gives you the interpretation.” The challenge of interpreting licenses intensifies when considering software that crosses borders and licenses that were written in other legal and cultural contexts, as identified by $RI_{7}$: “[W]hen you are reading the laws sometimes, you have to interpret and it’s very hard to make sure that your interpretation is the right one. And since we are talking in an international context, sometimes it depends $\dots$ there [are] some cultural aspects that make you interpret[ ] the law in [different] ways. But when you are talking about something which is worldwide, it’s very difficult to interpret how the people that designed the rules thought about it.”
@@
Beyond this, some interviewees expressed a need for tooling support for some of the more difficult compliance tasks, as defined in Figure 4. For example, $RI_{7}$ noted that they did not know of any tools for verifying license compliance.
@@
Finding 13: Respondents believed that licensing tools should support some of the more difficult licensing tasks identified by developers and assist developers with communicating with legal experts for those tasks.
@@
Finding 14: Developers expressed mixed feelings on the potential of generative AI with respect to software licensing: the impacts it might have are not yet clear, and further study is required.
@@
Overall, licensing tools with sufficient capabilities may improve the process of performing licensing tasks in various ways, as long as they are sufficiently reliable and accessible.
Finding 15: Respondents conveyed that software licensing tools should be reliable and easily accessible by developers, such as through existing tools and platforms.
5. Discussion
@@ 5.1. Who is responsible for license compliance?
In Section 4.2.1, we observed that respondents were divided about who exactly should be responsible for handling license compliance tasks, but the most common answer was that legal experts should ultimately be responsible. Accordingly, in Section 4.2.2, we observed that when faced with licensing issues, respondents’ most-reported strategy was indeed to escalate to legal experts. However, we also observed that such experts are not always available to developers.
This result is in contrast to the findings of another recent work that considered how legal practitioners approach software license compliance challenges (Wintersgill et al., 2024), which found that legal professionals believed that compliance should be a team effort, but that the primary responsibility for compliance should fall on developers, who are most familiar with the project. Legal professionals may have limited technical training and are not closely involved in the code base. In their view, it was the role of legal teams to provide developers with the tools, resources, and training they need to succeed and to provide guidance on only the more challenging issues.
It would seem that both developers and legal practitioners believe that the ultimate responsibility for compliance should fall on the other party, and both provide valid arguments for why this should be the case. In truth, the answer likely lies somewhere in the middle, with an appropriate collaboration of legal and development teams that can bring the optimal mix of technical knowledge and legal background. However, as noted by Wintersgill et al. (Wintersgill et al., 2024) and Section 4.3.3, this balance can be difficult to strike, since legal experts and software developers may have different interests and goals. More research is needed to explore the interaction between roles and determine ways to facilitate effective collaboration.
@@ 5.2. Where to focus future efforts?
While our results show that licensing tasks may be infrequent for developers, their importance is undiminished: failing to comply with licenses can result in significant costs (Claburn, 2022; Vincent, 2022; Claburn, 2023; Gunningham et al., 2004; Vance, 2010; Gross, 2007), justifying focused investment in tools to assist with licensing tasks.
As shown in Section 4.1.1 and Section 4.1.2, developers perform software licensing tasks at different rates, and some tasks are more difficult than others. In particular, while verifying license compliance was reported as one of the development tasks developers perform most frequently, it was also cited as the most difficult task by far. Despite this, respondents were broadly unaware of tools capable of assisting with this task. Section 4.3.1 showed low tool adoption, mainly because of the lack of tooling better integrated into the software development process. As suggested in Section 4.3.3, licensing tools built into an IDE or into a commonly-used service like GitHub would make their capabilities easier to access for developers, while also mitigating usability concerns highlighted in Section 4.3.2. Increased accessibility may also mitigate the last minute timing of resolving licensing issues, as identified in Section 4.2.2.
The broad array of licensing factors considered (Section 4.1.3), roles involved (Section 4.2.1) and potential solutions to issues (Section 4.2.2) indicate that the licensing tools must be able to take into account the context provided by the user and adapt to the specific circumstances of a given project.
Section 4.3.3 also articulated the developers’ varying views on generative AI as it relates to software licensing. Our interviews revealed that while some are hopeful that AI-based tools could assist with software licensing in new, powerful ways, others expressed caution regarding their use for such a critical purpose. Therefore, we propose that an investigation into the suitability of generative AI for licensing tasks is warranted. An AI-based licensing tool would need to be capable of providing meaningful analyses of software licenses and software components’ licensing status, while also being reliable enough to avoid making inappropriate recommendations or providing services best left to legal practitioners.
@@ 6. Threats to Validity
Construct Validity threats concern the relationship between theory and observation. As with survey studies generally, a possible threat may be related to the fact that our study can only report participants’ perception, rather than actual practices. Interviews partially mitigated this threat as they allowed us to seek clarification and expansion of responses. However, different kinds of studies, e.g., ethnographic studies, would be required to better understand developers’ practices in the field.
Internal Validity threats concern factors internal to the study that could influence our findings. One possible threat is subjectiveness in coding interviews’ open-ended answers and interview transcripts. To mitigate such a bias, we utilized an open coding methodology with a procedure to discuss and resolve conflicts. We employed diverse strategies to locate participants (professional networks, Discord servers, etc.) in an attempt to increase the pool of different perspectives and minimize potential bias, but we are aware of the issues that may arise from low response rates and self-selection bias. We followed best practices in formulating survey and interview questions, making sure that questions were written clearly and concisely to avoid confusion and biased language and conducting a small pilot study. Interviews were limited to at most half an hour, meaning that not all of a participant’s views were likely heard. We limited confirmation bias in qualitative analysis by independent coding, discussing disagreements, and arriving at a consensus based on the data.
External Validity threats concern the generalizability of our findings. The conclusions drawn in this study apply only to the population that participated in our survey and follow-up interviews. Our results cannot be generalized to the larger population, but in light of the themes that emerged, we believe that other software developers likely share some of the same experiences and challenges. That said, our goal was not to claim generalizability but to identify potential pain points that hinder developers from achieving efficient and effective license compliance for their projects.
@@ 7. Conclusion
We present a comprehensive study of current software licensing challenges faced by developers through a survey of 58 developers and seven follow-up interviews, deriving ???? key findings encompassing: (i) how frequently different licensing tasks are performed; (ii) their difficulty; (iii) the roles developers perceive as addressing licensing issues, (iv) processes for resolving those issues; and (v) the benefits and shortcomings of licensing tools, and opportunities to improve them. These results reveal actionable insights for the creators of licensing tools by highlighting areas that need improvement, such as low perceived interactivity and shallow analyses, as well as identifying both common and difficult licensing tasks as areas that can most benefit from future license tool development.
Respondents had different strategies for handling licensing issues and differing views on who is responsible for addressing them. Although many reported performing licensing tasks infrequently and without licensing tools, they also highlighted difficult tasks, particularly verifying license compliance. As such, we recommend that future research and tool development focus on pain points and investigate the suitability of generative AI for licensing tasks.

### v2

@@ 4.2.3. Understanding licenses
In interviews, we sought to clarify how developers acquired an understanding of software licenses. $RI_{3}$, who had indicated that they had received informal license training (see Section 3.4), described a “bootcamp” for licensing consisting of “three sessions of 45 minutes.” Some interviewees described how they had gained an understanding of common licenses over the years, such as $RI_{5}$, who noted, “I generally know that if it’s MIT or Apache, I’m good to go … and those are the ones that I encounter most regularly.” $RI_{6}$ had a similar familiarity with their most-used licenses but acknowledged that this did not extend to licenses they had less experience with: “I’m familiar with … the [Creative Commons] sharealike, the MIT license…. But I wouldn’t say that I’m comfortable knowing what all the most common ones mean. I know the ones that I know because they are the ones that I use.” However, such an approach warrants a warning: as we have identified in our previous work (Wintersgill et al., 2024), developers sometimes believe themselves to acquire a “feel” for licenses in this way despite misunderstanding their terms in some capacity, resulting in a false sense of confidence that can be a source of conflict with legal experts.
[…]
Finding 8: Developers report acquiring an understanding of licenses they encounter primarily through experience, despite the risk of misunderstanding them. To comprehend unfamiliar licenses, they often turn to online resources, including Generative AI tools and web search.
@@ 4.3. RQ3: Licensing Tools
To understand if and how developers use tools to assist with software licensing tasks, we asked participants to describe their use of licensing tools. We refer to a “licensing tool” as any software that can facilitate licensing tasks through automated analysis or by providing additional (licensing) context or information.
@@ 4.3.1. Tool adoption
The interview participants elaborated, suggesting that the frequency with which licensing tasks are performed may be a factor in low tool use. $RI_{5}$ stated, “I’m not aware of tools [because of] the fact that [licensing is] not something that I have to deal with every day…” Similarly, $RI_{6}$ was unfamiliar with tools but mused that someone who engaged with licensing more might be more familiar: “I am less aware [of licensing tools] … [but] I would be completely unsurprised if there are established open source tools to assist in this [that] I am not personally aware of. But the people that work for me and do this when it is an important task very well may [know of them].”
Finding 9: Adoption of software licensing tools was low among respondents, who suggested that this may be due to the relatively low frequency at which licensing tasks are performed.
@@ 4.3.2. Tool benefits and shortcomings
Tools can also help respondents choose licenses to assign to their software. In their survey response, $RI_{7}$ indicated that they benefited from tools that could give “an overview of all licensing options [and support] the selection of the right option.” They also noted a capability of “[g]enerating terms that I can copy/paste,” though it is not clear whether this referred to providing the text of existing licenses or creating new licenses. In interviews, $RI_{5}$ and $RI_{7}$ both referred to using choosealicense.com (GitHub, Inc., ) to assist them when assigning licenses to software, praising its ability to “answer your questions and then recommend the best license,” as $RI_{7}$ put it.
We also asked the respondents to identify the difficulties or shortcomings that they had encountered with the tools. Two respondents indicated that they had encountered issues with licensing tools being cumbersome or difficult to use, with $R_{43}$ describing usability as “poor.” Others noted that tools were not able to account for user feedback: in their survey response, $RI_{7}$ called current tooling “[n]ot interactive enough,” and noted that it could be improved with an “[i]nteractive decision tree” or a “[r]eliable chat bot.” Other concerns included the clarity of tools’ outputs, such as $R_{43}$’s observation that the “[a]ctual effect of a license is obscure[d] to developers,” and $R_{37}$’s opinion that tools “[m]ay abstract away finer details.”
Finding 10: Developers appreciated tools’ help in identifying licenses, highlighting dependencies which may cause licensing issues, and choosing appropriate licenses for their software. However, tools are hindered by low usability, low interactivity, and shallow, unclear analyses.
@@ 4.3.3. Areas requiring better tool support
Given the low reported tool adoption, we investigated in interviews which licensing tasks respondents believed could most benefit from better tool support.
[…]
@@
Beyond this, some interviewees expressed a need for tooling support for some of the more difficult compliance tasks, as defined in Figure 4. For example, $RI_{7}$ noted that they did not know of any tools to verify license compliance.
@@
Finding 11: Developers expressed a desire for increased tool support to explain licenses and their terms, compare licenses, verify license compliance, and provide guidance on information to share with legal experts.
@@
Finding 12: Developers expressed mixed opinions on the potential of generative AI with respect to software licensing: the impacts it might have are not yet clear, and further investigation is required.
@@
In general, licensing tools with sufficient capabilities can improve the process of performing licensing tasks in various ways, as long as they are sufficiently reliable and accessible.
Finding 13: Respondents emphasized that licensing tools should be reliable and readily accessible, particularly through seamless integration into existing development tools and platforms (e.g., IDEs and GitHub).
5. Discussion and Implications
@@ 5.1. Who is responsible for license compliance?
In Section 4.2.1, we observed that respondents were divided about who exactly should be responsible for handling license compliance tasks, but the most common answer was that legal experts should ultimately be responsible. Consequently, in Section 4.2.2, we observed that when faced with licensing issues, respondents’ most-reported strategy was indeed to escalate to legal experts. However, we also note that such experts are not always available to developers.
This result is in contrast to the findings of our recent study regarding considered how legal practitioners approach software license compliance tasks (Wintersgill et al., 2024), which found that legal professionals believed that compliance should be a team effort, but that the primary responsibility for compliance should fall on developers, who are most familiar with the project. Legal professionals may have limited technical training and are not closely involved in the code base. In their view, it was the role of legal teams to provide developers with the tools, resources, and training they need to succeed and to provide guidance on only the more challenging issues.
Both developers and legal practitioners appear to believe that the ultimate responsibility for compliance should fall on the other party, and both provide valid arguments for why this should be the case. In reality, the answer likely lies somewhere in the middle, with an appropriate collaboration of legal and development teams that can bring the optimal combination of technical knowledge and legal background. However, as noted in our previous work (Wintersgill et al., 2024) and in Section 4.3.3, this balance can be difficult to strike, since legal experts and software developers may have different interests and goals. More research is needed to explore the interaction between roles and determine ways to facilitate effective collaboration. This can involve understanding what materials legal experts need from developers to effectively analyze potential issues, while simultaneously providing developers with tools to package that information in a way that is accessible and understandable to legal practitioners.
@@ 5.2. Where to focus future efforts?
Although our results show that some licensing tasks may be infrequent for developers, their importance is undiminished: failure to comply with licenses can result in significant costs (Claburn, 2022; Vincent, 2022; Claburn, 2023; Gunningham et al., 2004; Vance, 2010; Gross, 2007), justifying focused investment in licensing tools.
As shown in Section 4.1.1 and Section 4.1.2, developers perform software licensing tasks at different rates, and some tasks are more difficult than others. While verifying license compliance was reported as one of the development tasks developers perform most frequently, it was also cited as the most difficult task by far. Despite this, respondents were broadly unaware of tools capable of assisting with this task. Section 4.3.1 showed low tool adoption, mainly due to the lack of tools better integrated into the software development process. As suggested in Section 4.3.3, licensing tools built into an IDE or a commonly-used service like GitHub would make them easier to access for developers while also mitigating usability concerns highlighted in Section 4.3.2. Increased accessibility may also mitigate the last-minute timing of resolving licensing issues, as identified in Section 4.2.2. The broad array of licensing factors considered (Section 4.1.3), roles involved (Section 4.2.1), and potential solutions to issues (Section 4.2.2) indicate that licensing tools must be able to take into account the context provided by the user and adapt to the circumstances of a given project.
Section 4.3.3 also articulated developers’ varying views on generative AI as it relates to software licensing. Our interviews revealed that while some are hopeful that AI-based tools can assist with software licensing in new and powerful ways, others expressed caution regarding their use for such a critical purpose. Therefore, we propose that an investigation into the suitability of generative AI for licensing tasks is warranted. An AI-based licensing tool would need to be capable of providing meaningful analyses of software licenses and software components’ licensing status while also being reliable enough to avoid making inappropriate recommendations or providing services best left to legal practitioners.
@@ 6. Threats to Validity
Construct Validity. As with survey studies in general, our study cn only report participants’ perceptions rather than actual practices. The interviews partially mitigated this threat, as they allowed participants to clarify and expand on the survey responses. However, studies such as ethnographic studies would be required to better understand developers’ practices in the field.
Internal Validity. An additional possible threat is subjectivity in coding interviews’ open-ended answers and transcripts. To mitigate such bias, we utilized an open-coding methodology with a rigorous procedure to discuss and resolve conflicts. We employed diverse strategies to locate participants (professional networks via email and LinkedIn, Discord servers, snowball sampling) in an attempt to expand the pool to encompass different perspectives and minimize potential bias. We followed best practices in formulating survey and interview questions, making sure that the questions were written clearly and concisely to avoid confusion and biased language, and conducted a small pilot study. To encourage and facilitate participation from developers online, the survey and interviews were designed to be completed quickly: the survey could be finished in about ten minutes, and interviews were limited to at most 30 minutes, meaning that not all of a participant’s views were likely heard. We limited confirmation bias in qualitative analysis by independent coding, discussing disagreements, and arriving at a consensus based on the data.
External Validity. Considering our study’s response rate and the potential for self-selection bias, the conclusions drawn in this study apply only to the population that participated in our survey and follow-up interviews. Our results cannot be generalized to the larger population, but in light of the themes that emerged, we believe that other software developers likely share some of the same experiences and challenges. That said, our goal is not to claim generalizability but to illustrate the current state of practice and to identify potential pain points that hinder developers from achieving efficient and effective license compliance for their projects.
@@ 7. Conclusion
We present a study of current software licensing challenges faced by developers through a survey of 58 developers and 7 follow-up interviews, deriving ???? key findings encompassing: (i) how frequently different licensing tasks are performed; (ii) their difficulty; (iii) the roles developers perceive as addressing licensing issues, (iv) processes for resolving those issues; and (v) the benefits and shortcomings of licensing tools. These results reveal actionable insights for the creators of licensing tools by highlighting areas that need improvement, such as low perceived interactivity and shallow analyses, as well as identifying both common and difficult licensing tasks as areas that can most benefit from future license tool development.
The respondents had different strategies for handling licensing issues and differing views on who is responsible for addressing them. Although many reported performing licensing tasks infrequently and without licensing tools, they also highlighted difficult tasks, particularly verifying license compliance. As such, we recommend that future research and tool development focus on pain points and investigate the suitability of generative AI for licensing tasks.